40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 211     Entry time: Sat Dec 22 00:52:57 2007
Author: tobin 
Type: Configuration 
Category: PSL 
Subject: ISS surgery 
In an attempt to quell oscillations in the (unused) outer loop portion of the ISS, I shorted the "PD+" and "PD-" signals from the (nonexistent) outer-loop diodes, and soldered in 47pf compensation capacitors in C92 and C220. This seems to have eliminated oscillations seen at TP41 and TP42. There's still something amiss at TP30 and maybe TP20. Otherwise, the ISS seems happy. I can turn the gain slider to +15dB without saturation (with the HEPA off), though there seems to be less light on the diode (~3.9V) than a week or two ago.
ELOG V3.1.3-